### v1

# LaVy: Vietnamese Multimodal Large Language Model

Abstract
Large Language Models (LLMs) and Multimodal Large language models (MLLMs) have taken the world by storm with impressive abilities in complex reasoning and linguistic comprehension. Meanwhile there are plethora of works related to Vietnamese Large Language Models, the lack of high-quality resources in multimodality limits the progress of Vietnamese MLLMs. In this paper, we pioneer in address this by introducing LaVy, a state-of-the-art Vietnamese MLLM, and we also introduce LaVy-Bench benchmark designated for evaluating MLLMs’s understanding on Vietnamese visual language tasks. All code and model weights are public at https://github.com/baochi0212/LaVy
1 Introduction
In recent years, Large Language Models (LLMs) have demonstrated remarkable capabilities in various natural language processing tasks, showcasing their proficiency in complex reasoning and linguistic comprehension. The success of LLMs has inspired researchers to explore the potential of Multimodal Large Language Models (MLLMs), which incorporate visual information alongside textual data. MLLMs have shown promising results in tasks that require understanding the interplay between language and vision, such as image captioning, visual question answering, and multimodal machine translation.
While there has been significant progress in developing Vietnamese LLMs, the lack of high-quality multimodal resources has hindered the advancement of Vietnamese MLLMs. The availability of diverse and well-annotated datasets is crucial for training and evaluating MLLMs, as they rely on the integration of visual and textual information to perform multimodal tasks effectively.
To address this limitation and foster research in Vietnamese multimodal language understanding, we introduce LaVy, Vietnamese first MLLM and achieve state-of-the-art performance in Vietnamese vision language tasks. LaVy is designed to leverage the rich visual and linguistic information present in Vietnamese data, enabling it to tackle a wide range of multimodal tasks with improved performance. Our model outperforms a multilingual baseline mBLIP Geigle et al. (2023) on different tasks by a large margin. By developing LaVy, we aim to bridge the gap between Vietnamese LLMs and MLLMs, providing researchers and practitioners with a powerful tool for exploring the intersection of language and vision in the Vietnamese context.
Furthermore, to facilitate the evaluation and comparison of Vietnamese MLLMs, we propose the LaVy-Bench benchmark. This benchmark consists an open VQA task and an in-the-wild test set, specifically designed to assess the visual language understanding and generation capabilities of MLLMs in the Vietnamese and in-the-wild images. By establishing a standardized evaluation framework, we aim to promote the development and benchmarking of Vietnamese MLLMs, driving innovation and collaboration within the research community.
In this paper, we present LaVy and the LaVy-Bench benchmark as significant contributions to the field of Vietnamese multimodal language understanding. We provide a detailed description of LaVy’s architecture, data curation and training procedure. Additionally, we introduce the LaVy-Bench benchmark, discussing its design principles, task composition, and evaluation metrics. Through extensive experiments and analysis, we demonstrate the effectiveness of LaVy and the utility of the LaVy-Bench benchmark in advancing Vietnamese MLLM research.
2 Related work
2.1 Large Language Model
Recent advancements in Large Language Models (LLMs) have showcased remarkable capabilities across various natural language processing tasks, including dialogue, creative writing, and problem-solving. Models such as LLaMA Touvron et al. (2023a, b), Mistral Jiang et al. (2023), and Gemma Mesnard et al. (2024) have leveraged scalable Transformer-based architectures Vaswani et al. (2017) and large-scale data to become foundation models for general reasoning tasks. These models have demonstrated impressive performance and have set new benchmarks in the field.
Following the trend of LLMs, several Vietnamese language models emerged such as PhoGPT Nguyen et al. (2023b), Vistral Nguyen et al. (2023a) perform outstandingly in Vietnamese LLM benchmarks and NLP tasks.
2.2 Multimodal Large Language Model
Witness the exceptional performance of GPT-4 Achiam et al. (2023) and Gemini Pro Vision Anil et al. (2023) in visual language tasks, recent research has focused on developing Multimodal Large Language Models (MLLMs) to achieve unified understanding and reasoning across different modalities, building upon the success of Large Language Models (LLMs). Various methods have been proposed to integrate information from multiple modalities into pre-trained LLM architectures. For instance, Flamingo Alayrac et al. (2022) and BLIP-2 Li et al. (2023) introduce different techniques for fusing visual tokens with frozen LLMs using gated attention or Q-former. Inspired by the effectiveness of instruction tuning, LLaVA Liu et al. (2023) and MiniGPT-4 Zhu et al. (2024) align visual input with LLMs through visual instruction tuning, demonstrating impressive results. Another active line of work is researching efficient MLLMs, resulting in lightweight model families like Bunny He et al. (2024). Meanwhile, recent works are pioneering development in vision-language tasks for low-resource languages, such as Peacock Alwajih et al. (2024)
3 LaVy
3.1 Architecture
Our model are built with LlaVA architecture Liu et al. (2023) using three main components:
  • Vision Encoder: The CLIP-Large model, developed by Radford et al. (2021), is used as the vision encoder.
  • MLP Projector: A two-layer Multi-Layer Perceptron (MLP) projector is employed to align the output representations from the visual and language modalities. This projector ensures that the visual and textual information is transformed into a common space.
  • Language Model: The third component is a Large Language Model, which is a language model responsible for generating textual information, take the aligned representations from the MLP projector.
3.2 Data Curation
The barrier for Vietnamese MLLMs’ development is resource for training, which is handled by the our novel pipeline of data collection.
  • Translated then Refined: We utilize LlaVA training data composing of filtered 558K LAION-CC-SBU captions and 150K GPT-generated multimodal instructions. With the acknowledgement of the insufficient competency of open-source translation projects and translation API, we firstly translate a sample with VinAI translation, then we prompt Gemini Pro to rewrite it in more accurate and natural way for Vietnamese language by pairing translated sample and original English sample in Gemini prompt. Because the captions dataset is massive, we only refine a subset of 150K randomly sampled captions, and combine it with 558K non-write captions to finalize a captioning dataset. Finally, GPT-generated instructions is all refined to make 150K high-quality Vietnamese instructions
[…]
In eventuality, we curated Vietnamese datasets with 708K image-caption pairs for pretraining and 166K high-quality instructions for finetuning in training step. Our pipeline is clearly illustrated in Image 1.
Image 1: Data pipeline
3.3 Training Procedure
The training procedure is divided into 2 steps:
  • Pretraining: Align the vision embeddings from a pre-trained vision encoder with the text embeddings from the LLM by optimizing only the cross-modality projector using a cross-entropy loss for next token prediction.
  • Finetuning: We apply visual instruction tuning to fully utilize the MLLM’s capabilities across different multimodal tasks. We use the same cross-entropy loss as in the pre-training stage, but this time, they employ Low-Rank Adaptation (LoRA) to train both the cross-modality projector and the LLM backbone.
4 Experiment
4.1 Implementation details
We use Vistral 7B as LLM backbone and CLIP large visual encoder. The training process for the LaVy consists of two stages. In the first stage, the model is pretrained using a dataset of 708k captions for 1 epoch, with a global batch size of 64 and a learning rate of 1e-3. During this stage, all model parameters are frozen, except the MLP layers. Besides, we don’t shuffle data but train the model to learn from unrefined data to refined data.
The second stage involves finetuning the model using an instructions dataset. This stage also spans 1 epoch, with a global batch size of 32 and a learning rate of 2e-5. In this phase, only the newly introduced LoRA (Low-Rank Adaptation) parameters are trainable.
Besides, in evaluation, we apply greedy decoding to generate all models’ responses Lin and Chen (2023)
4.2 LaVy-Bench
We construct LaVy-Bench to benchmark models’ Vietnamese Visual Language understanding.
4.2.1 Zero-shot Visual Question Answering (VQA)
We evaluate the zero-shot Visual Question Answering (VQA) performance of models on the OpenViVQA Nguyen et al. (2023c) dev set, which consists of 3,505 samples. This dataset challenges the models’ understanding of the relationships between Vietnamese images and natural language. Furthermore, we propose a new metric for automatic evaluation to replace older metrics, such as BLEU Papineni et al. (2002), which do not accurately reflect models’ competency in the VQA tasks. Our metric is inspired by LLM-as-a-Judge Zheng et al. (2023), which utilizes Gemini Pro to verify the accuracy of generated responses for question-answer pairs. In Table 1, it’s apparent that LaVy’s zero-shot VQA performance (33.5%) outshadows mBLIP-Bloomz-7B (27.9%) and mBLIP-mT0-XL-5B (20.0%). However, the OpenViVQA dataset contains a variety of question types that do not appear in our training dataset, such as OCR, making this test set particularly challenging for our model, not mention to our traning instructions just includes descriptions of 8.000 Vietnamese crawled images.
Model Accuracy
mBLIP (mT0-XL-5B) 20.0
mBLIP (BLOOMZ-7B) 27.9
LaVy 33.5
Table 1: Zero-shot VQA on OpenViVQA dev set. Models’ output accuracy are evaluated by Gemini Pro
4.2.2 In-the-wild benchmark
To further assess the models’ comprehension, we follow the evaluation methodology LLaVA benchmark (in-the-wild) Liu et al. (2023) and recollect set of 24 diverse images and 60 questions in 3 main types: Complex Reasoning, Detail Description and Conversation. The collected images and manually crafted questions aim to diversify the test set in various aspects: culture, race, image types,… and avoiding opting for images in crawled training images. For each question, we then prompt Gemini Pro Vision to generate detailed description of image and answer the question, and finally use Gemini Pro to rate the models’ response on the scale of 1 to 5 while taking Gemini Pro’s response as ground-truth. We opted against using detailed descriptions as references (unlike LlaVA) due to their occasional irrelevance. Instead, we found Gemini Pro Vision’s in-the-wild benchmark responses to be closely pertinent. Then we ask model to score outputs in 3 criteria: relevancy, accuracy and naturalness, and sum them all for final score at the end
In comparison with mBLIP baselines in Table 2, LaVy outperforms sharply in all types of questions: Conversation (+30%), Detail Description (+64%) and Complex Reasoning (+49%). In overall, our model is scored 60.6% by Gemini Pro. Some qualitative test cases are depicted in Table 3.
Model Conversation Detail description Complex reasoning All
mBLIP (mT0-XL-5B) 48.8 26.5 28.3 34.5
mBLIP (BLOOMZ-7B) 53.0 28.6 44.1 41.9
LaVy 69.0 46.9 65.9 60.6
Table 2: Performance on in-the-wild benchmark
5 Limitations
Our model still have several limitations:
  • Although LaVy exhibits deep understanding in Vietnamese visual language tasks but still faces many challenges for example: OCR, Object Counting, … due to lack of high-quality annotated data for these tasks.
  • Moreover, like other MLLMs, our model still suffers from hallucination where it generates irrelevant information, redundant details or misinformation.
6 Conclusion
In this paper, we have introduced LaVy, a pioneering state-of-the-art Vietnamese Multimodal Large Language Model (MLLM) that aims to address the lack of high-quality resources in multimodality for the Vietnamese language. LaVy represents a significant step forward in the development of Vietnamese MLLMs, enabling complex reasoning and linguistic comprehension in tasks that involve both visual and textual information.
Furthermore, we have presented LaVy-Bench, a comprehensive benchmark designed specifically for evaluating the performance of MLLMs on Vietnamese visual language tasks. This benchmark provides a standardized platform for assessing the capabilities of Vietnamese MLLMs, facilitating the comparison and advancement of these models. Our model also have proved SOTA perfomance in comparison with mBLIP baselines in test sets of benchmark.
As future work, we plan to expand the capabilities of LaVy by incorporate diverse instructions to entirely handle challenging tasks like Vietnamese OCR, Object Counting. We hope our work will contribute to the advancement of Vietnamese MLLMs’ development.